# Conditioning LLMs to Generate Code-Switched Text

### v2

@@ 6 Automatic Evaluation
6.1 Reference-based metrics
Model BLEU BERTScore chrF
Llama3 8B 34.49 81.64 53.17
Llama3 8B Instruct 33.42 81.77 52.01
Mistral 31.65 80.93 50.56
Mistral Instruct 25.98 78.66 44.58
Llama3.3-70Bfs 22.41 79.77 44.57
GPT-4ofs 32.25 83.09 50.48
NLLB 35.56 84.11 54.74
Identity 33.34 82.31 45.51
@@ 6.1 Reference-based metrics
We report the results of BLEU (Papineni et al., 2002), BERTScore (Zhang et al., 2020), and chrF (Popović, 2015), implemented with the evaluate library. All three are task-agnostic quality metrics that give results between 0-1, based on character-level F-score, n-gram precision and semantic similarity using contextual embeddings [^4] BERTscore has been calculated using the embeddings from the model Bert Base Multilingual Cased. respectively. We compute the metrics for all systems, and include an Identity system that simply returns the provided input as the output.
@@
The results of the evaluation can be seen in Table 5. The best model is NLLB, with the highest scores for the three metrics. It is closely followed by GPT-4ofs, with the second highest BERTScore, and fine-tuned Llama3, with the second highest BLEU and chrF. They are closely followed by Llama3 Instruct and Mistral. The Identity system scores nearly as well as the top-performing models. The strong results from the Identity baseline and few-shot models, which often produce monolinugal outputs, suggest that reference-based metrics assign high scores to models that match only the English part of the reference. This reflects the nature of the task and dataset, and highlights the limitations and artifacts of using reference-based metrics to evaluate code-switched generation.
6.2 GPT as a judge
@@
As a complementary automatic assessment of the outputs of our models, we have implemented a zero-shot pair-wise evaluation using GPT-4o as a judge, mimicking the settings of the human evaluation. Details about the implementation are included in Appendix G. Results are shown in the third and fourth columns of Table 4. GPT shows a strong preference for few-shot models, whereas these models are ranked third- and second-to-last by humans. Based on the error analysis (c.f Section 5.2), few-shot models tend to make many CS-related mistakes, although they are the most fluent. Thus, this disagreement may be caused by the humans adhering to the guidelines and taking the presence of CS as the main criterion, whereas GPT is making decisions based on the style and fluency of the answers. Regarding fine-tuned LLMs, they all share the same relative ranking in both evaluations. Finally, NLLB is ranked last by GPT, while it is the fourth best model overall as ranked by humans. Further research is needed to explain this behaviour, but there may be some stylistic features of the NLLB model’s outputs that are affecting GPT’s preferences.
6.3 Correlation With Human Evaluation
@@
The reference-based metrics used in Section 6.1 are known to have weak correlations with human judgment in NLG tasks (Sai et al., 2022), whereas JudgeLLM-based evaluations seem promising (Chiang and Lee, 2023; Wang et al., 2023). In this section, we compare reference-based metrics and GPT scores with the preference-based scores obtained in Section 5.1.
We calculate Pearson’s ($\rho$) correlation coefficient at instance-level, using the $700$ instances employed for the error classification and human evaluation (the output of 7 models for $100$ source sentences). [^5] We do not consider the reference CS sentences when calculating the correlations. Each data point corresponds to the CS output of one particular model for an English source sentence, and we compute the correlation using two values: the score obtained by the model for this instance in the human preference-based evaluation of Section 5.1, and the score it attains if we apply the same strategy using the values of the reference-based metrics to determine the winner, or, in case of JudgeLLM, the scores given by GPT.
@@
Figure 2: Heatmap of the correlations between human scores and reference-based metrics and scores given by GPT, calculated using the Pearson Correlation Coefficient. The correlations are calculated for all instances, as well as for different subsets of instances, according to the type of errors they exhibit.
@@
The correlation coefficients are shown in Figure 2. The top part of the figure shows the correlation using all the instances, whereas the bottom part only considers those instances that showed some type of error, according to the error analysis described in Section 5.2.
If we consider all the instances, the maximum $\rho$ correlation value with reference-based metrics is $0.22$, which indicates a low alignment with human scores. The metric with highest correlation is chrF, in line with previous research (Popović, 2015). GPT-4ofs shows a $\rho$ of 0.53, which is stronger than reference-based metrics, but still too weak to be regarded as a reliable measure for assessing CS generation.
@@
If we instead consider instances with errors from the human evaluation, there is again a higher correlation between human scores and GPT’s judgements, with a margin of at least 0.28 points. Instances with CS errors show the lowest overall correlation. This likely derives from the fact that human evaluators never prefer an instance without CS as instructed in the guidelines, but reference-based and JudgeLLM-based metrics are not sensitive to these nuances, and may assign high scores to instances regardless of whether they contain CS or not.
@@
All in all, these results confirm that several of the most commonly used reference-based metrics for NLG have a weak correlation with human judgments when evaluating CS generation. This underscores the need to research more specialized evaluation methods designed specifically to capture the nuances of this task that correlate with human judgement.
7 Conclusion
@@
We only perform an in-domain evaluation where the train, validation and test sets had the same origin. Additionally, we would like to test the efficiency of our models in an out-of-domain setting, since one of the use-cases of a CS generation model is to create parallel corpora to evaluate the abilities of models to perform different tasks when there is CS.
@@
Finally, as we have pointed out, we are aware of the problems of the automatic metrics that we have used to evaluate the outputs of our models, that do not capture the nuances of our task. In the future, we would like to investigate how to improve this evaluation by designing new methods to automatically evaluate CS generation, focusing on a more linguistic approach able to capture the linguistic and social intricacies of CS.

### v3

@@ 6.  Automatic Evaluation
6.1.  Reference-based metrics
@@
We report the results of BLEU (Papineni et al., 2002), BERTScore (Zhang et al., 2020), and chrF (Popović, 2015).  [^5] The metrics are implemented using the evaluate library. All three are task-agnostic quality metrics that give results between 0-1, based on character-level F-score, n-gram precision and semantic similarity using contextual embeddings [^6] BERTscore has been calculated using the embeddings from the model Bert Base Multilingual Cased. respectively. We compute the metrics for all systems, and include an Identity system that simply returns the provided input as the output.
@@
The results of the evaluation can be seen in Table 5. In the in-domain setting, the best model is NLLB, with the highest scores for the three metrics. It is closely followed by GPT-4ofs, with the second highest BERTScore, and fine-tuned Llama3, with the second highest BLEU and chrF. Llama3 Instruct follows closely. The Identity system scores nearly as well as the top-performing models for the in-domain set and obtains the best scores out-of-domain. The strong results from the Identity baseline and few-shot models, which often produce monolingual outputs, suggest that reference-based metrics assign high scores to models that match only the English part of the reference. This reflects the nature of the task and dataset, and highlights the limitations and artifacts of using reference-based metrics to evaluate code-switched generation. Comparing these results with the out-of-domain evaluation, we observe that all models achieve considerably lower scores. Although GPT-4ofs obtains the highest scores, the margin is relatively small. This drop in performance is likely due to specific characteristics of the out-of-domain dataset that negatively affect all models’ predictions, such as longer sentence lengths.
6.2.  GPT as a judge
As a complementary automatic assessment of the outputs of our models, we have implemented a zero-shot pair-wise evaluation using GPT-4o as a judge, mimicking the settings of the human evaluation. Details about the implementation are included in Appendix G. Results are shown in Table 6.
GPT
Model Score Rank
Gold Standard 744.5 [Uncaptioned image]
GPT-4ofs 603.0 [Uncaptioned image]
Llama3.3-70Bfs 582.0 [Uncaptioned image]
Llama3 8B 440.5 4
Llama3 8B Instruct 410.0 5
NLLB 370.0 6
Table 6: Ranking of models using GPT as a judge.
@@ 6.2.  GPT as a judge
GPT shows a strong preference for few-shot models, whereas these models are ranked last and second-to-last by humans. Based on the error analysis (c.f Section 5.2), few-shot models tend to make many CS-related mistakes, as they often output purely monolingual sentences, although they are the most fluent overall. This suggests that this disagreement may be caused by the humans adhering to the guidelines and taking the presence of CS as the main criterion, whereas GPT is making decisions based on the style and fluency of the answers. Regarding fine-tuned LLMs, they share the same relative ranking in both evaluations, with Llama3 being the preferred model by humans and GPT. Finally, NLLB is ranked last by GPT, while it is the second best model overall as ranked by humans. Further research is needed to explain this behaviour, but there may be some stylistic features of the NLLB model’s outputs that are affecting GPT’s preferences.
6.3.  Correlation With Human Evaluation
@@
The reference-based metrics used in Section 6.1 are known to have weak correlations with human judgment in NLG tasks (Sai et al., 2022), whereas JudgeLLM-based evaluations seem promising (Chiang and Lee, 2023; Wang et al., 2023). In this section, we compare reference-based metrics and GPT scores with the preference-based scores obtained in Section 5.1, allowing us to examine how closely these automatic metrics align with human preferences.
@@
Figure 2: Heatmap of the correlations between human scores and reference-based metrics and scores given by GPT, calculated using the Pearson Correlation Coefficient. The correlations are calculated for all instances, as well as for different subsets of instances, according to the type of errors they exhibit. * indicates statistical significance $(p\leq 0.05)$.
We calculate Pearson’s ($\rho$) correlation coefficient at instance-level, using the $1,000$ instances employed for the error classification and human evaluation (the output of $5$ models for $200$ source sentences). [^7] We do not consider the reference CS sentences when calculating the correlations. Each data point corresponds to the CS output of one particular model for an English source sentence, and we compute the correlation using two values: the score obtained by the model for this instance in the human preference-based evaluation of Section 5.1, and the score it attains if we apply the same strategy using the values of the reference-based metrics to determine the winner, or, in case of JudgeLLM, the scores given by GPT. This allows us to directly compare the two types of evaluations on an instance-by-instance basis.
@@
The correlation coefficients are shown in Figure 2. The top part of the figure shows the correlation using all the instances, whereas the bottom part only considers those instances that showed some type of error, according to the error analysis described in Section 5.2. These results provide a more detailed view of how the metrics behave across different subsets of instances. If we consider all the instances, the maximum $\rho$ correlation value with reference-based metrics is $0.09$, which indicates a low alignment with human scores. GPT-4ofs shows a $\rho$ of $0.35$, which is stronger than reference-based metrics, but still too weak to be regarded as a reliable measure for assessing CS generation.
@@
If we instead consider instances with errors from the human evaluation, there is again a higher correlation between human scores and GPT’s judgments, with a margin of at least $0.25$ points. Instances with CS errors show the lowest overall correlation with reference-based metrics. This likely derives from the fact that human evaluators never prefer an instance without CS as instructed in the guidelines, but reference-based metrics are not sensitive to these nuances, and may assign high scores to instances regardless of whether they contain CS or not. GPT, however, obtains the highest correlation in these kinds of instances.
@@
All in all, these results confirm that several of the most commonly used reference-based metrics for NLG have a weak correlation with human judgments when evaluating CS generation. This underscores the need to research more specialized evaluation methods designed specifically to capture the nuances of this task that correlate with human judgment.
7.  Conclusion
@@
Finally, as we have pointed out, we are aware of the problems of the automatic metrics that we have used to evaluate the outputs of our models, which do not capture the nuances of our task. In the future, we would like to investigate how to improve this evaluation by designing new methods to automatically evaluate CS generation, focusing on a more linguistic approach able to capture the linguistic and social intricacies of CS.
Acknowledgments
This work is supported by the projects DeepMinor (CNS2023-144375) funded by MTDFP/ and by European Union Next GenerationEU/ PRTR; HumanAIze (AIA2025-163322-C61) funded by MICIU/AEI/10.13039/501100011033; and DeepThought (PID2024-159202OB-C21) funded by MICIU/AEI /10.13039/501100011033 and by ERDF, EU. Maite Heredia holds a PhD grant from the University of the Basque Country UPV/EHU (PIF23/218).
Bibliographical References
  • G. Aguilar, S. Kar, and T. Solorio (2020) LinCE: a centralized benchmark for linguistic code-switching evaluation. In Proceedings of the Twelfth Language Resources and Evaluation Conference, N. Calzolari, F. Béchet, P. Blache, K. Choukri, C. Cieri, T. Declerck, S. Goggi, H. Isahara, B. Maegaard, J. Mariani, H. Mazo, A. Moreno, J. Odijk, and S. Piperidis (Eds.), Marseille, France, pp. 1803–1813 (English). External Links: Link, ISBN 979-10-95546-34-4 Cited by: §1, §2, §3.1, §3.
  • R. Appel and P. C. Muysken (2005) Language contact and bilingualism. Amsterdam University Press. Cited by: §1.
  • J. Barnes, N. Perez, A. Bonet-Jover, and B. Altuna (2025) Summarization metrics for Spanish and Basque: Do automatic scores and LLM-judges correlate with humans?. External Links: 2503.17039, Link Cited by: §3.2.1.
  • R. Bawden and F. Yvon (2023) Investigating the translation performance of a large multilingual language model: the case of BLOOM. In Proceedings of the 24th Annual Conference of the European Association for Machine Translation, M. Nurminen, J. Brenner, M. Koponen, S. Latomaa, M. Mikhailov, F. Schierl, T. Ranasinghe, E. Vanmassenhove, S. A. Vidal, N. Aranberri, M. Nunziatini, C. P. Escartín, M. Forcada, M. Popovic, C. Scarton, and H. Moniz (Eds.), Tampere, Finland, pp. 157–170. External Links: Link Cited by: §4.
  • C. Callison-Burch, C. Fordyce, P. Koehn, C. Monz, and J. Schroeder (2007) (Meta-) evaluation of machine translation. In Proceedings of the Second Workshop on Statistical Machine Translation, C. Callison-Burch, P. Koehn, C. S. Fordyce, and C. Monz (Eds.), Prague, Czech Republic, pp. 136–158. External Links: Link Cited by: §3.2.1.
  • C. Chiang and H. Lee (2023) Can large language models be an alternative to human evaluations?. In Proceedings of the 61st Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), A. Rogers, J. Boyd-Graber, and N. Okazaki (Eds.), Toronto, Canada, pp. 15607–15631. External Links: Link, Document Cited by: §6.3, §6.
  • A. Chowdhery, S. Narang, J. Devlin, M. Bosma, G. Mishra, A. Roberts, P. Barham, H. W. Chung, C. Sutton, S. Gehrmann, P. Schuh, K. Shi, S. Tsvyashchenko, J. Maynez, A. Rao, P. Barnes, Y. Tay, N. Shazeer, V. Prabhakaran, E. Reif, N. Du, B. Hutchinson, R. Pope, J. Bradbury, J. Austin, M. Isard, G. Gur-Ari, P. Yin, T. Duke, A. Levskaya, S. Ghemawat, S. Dev, H. Michalewski, X. Garcia, V. Misra, K. Robinson, L. Fedus, D. Zhou, D. Ippolito, D. Luan, H. Lim, B. Zoph, A. Spiridonov, R. Sepassi, D. Dohan, S. Agrawal, M. Omernick, A. M. Dai, T. S. Pillai, M. Pellat, A. Lewkowycz, E. Moreira, R. Child, O. Polozov, K. Lee, Z. Zhou, X. Wang, B. Saeta, M. Diaz, O. Firat, M. Catasta, J. Wei, K. Meier-Hellstern, D. Eck, J. Dean, S. Petrov, and N. Fiedel (2023) PaLM: scaling language modeling with pathways. Journal of Machine Learning Research 24 (240), pp. 1–113. External Links: Link Cited by: §1.
  • Cohere For AI (2024) C4ai-command-r-v01 (revision 8089a08). External Links: Link, Document Cited by: §3.2, §3.
  • T. Dettmers, A. Pagnoni, A. Holtzman, and L. Zettlemoyer (2023) QLoRA: efficient finetuning of quantized llms. External Links: 2305.14314, Link Cited by: §4.
  • J. Dewaele and L. Wei (2014) Attitudes towards code-switching among adult mono- and multilingual language users. Journal of Multilingual and Multicultural Development 35 (3), pp. 235–251. External Links: Document, Link, https://doi.org/10.1080/01434632.2013.859687 Cited by: §2.
  • A. Dubey, A. Jauhri, A. Pandey, A. Kadian, A. Al-Dahle, A. Letman, A. Mathur, A. Schelten, A. Yang, A. Fan, et al. (2024) The llama 3 herd of models. arXiv preprint arXiv:2407.21783. Cited by: §4.
  • K. A. Dwyer (2017) My brother the mexican. PORTAL Journal of Multidisciplinary International Studies 14, pp. . External Links: Document Cited by: §4.
  • K. A. Dwyer (2019a) Doce horas: a family border tale. PORTAL Journal of Multidisciplinary International Studies 16, pp. 163–165. External Links: Document Cited by: §4.
  • K. A. Dwyer (2019b) Gringos mexicanos. PORTAL Journal of Multidisciplinary International Studies 16, pp. 160–162. External Links: Document Cited by: §4.
  • K. A. Dwyer (2019c) La manda. PORTAL Journal of Multidisciplinary International Studies 16, pp. 156–159. External Links: Document Cited by: §4.
  • K. A. Dwyer (2019d) Simón. PORTAL Journal of Multidisciplinary International Studies 16, pp. 153–155. External Links: Document Cited by: §4.
  • K. A. Dwyer (2020) La vaca y coatlicue. PORTAL Journal of Multidisciplinary International Studies 17, pp. 179–181. External Links: Document Cited by: §4.
  • T. Fu, D. Cai, L. Liu, S. Shi, and R. Yan (2024) Disperse-then-merge: pushing the limits of instruction tuning via alignment tax reduction. In Findings of the Association for Computational Linguistics: ACL 2024, L. Ku, A. Martins, and V. Srikumar (Eds.), Bangkok, Thailand, pp. 2967–2985. External Links: Link, Document Cited by: §5.2.
  • B. Gregorius and T. Okadome (2022) Generating code-switched text from monolingual text with dependency tree. In Proceedings of the 20th Annual Workshop of the Australasian Language Technology Association, P. Parameswaran, J. Biggs, and D. Powers (Eds.), Adelaide, Australia, pp. 90–97. External Links: Link Cited by: §2.
  • D. Gupta, A. Ekbal, and P. Bhattacharyya (2020) A semi-supervised approach to generate the code-mixed text using pre-trained encoder and transfer learning. In Findings of the Association for Computational Linguistics: EMNLP 2020, T. Cohn, Y. He, and Y. Liu (Eds.), Online, pp. 2267–2280. External Links: Link, Document Cited by: §2.
  • M. Heredia, M. Fernández Trinidad, and M. Jiménez-Bravo (2025) Actitudes lingüísticas hacia el cambio de código entre valenciano y castellano. Revista Española de Lingüística Aplicada/Spanish Journal of Applied Linguistics. External Links: ISSN 0213-2028, Link, Document Cited by: §2.
  • I. Hsu, A. Ray, S. Garg, N. Peng, and J. Huang (2023) Code-switched text synthesis in unseen language pairs. In Findings of the Association for Computational Linguistics: ACL 2023, A. Rogers, J. Boyd-Graber, and N. Okazaki (Eds.), Toronto, Canada, pp. 5137–5151. External Links: Link, Document Cited by: §2.
  • S. Khanuja, S. Dandapat, A. Srinivasan, S. Sitaram, and M. Choudhury (2020) GLUECoS: an evaluation benchmark for code-switched NLP. In Proceedings of the 58th Annual Meeting of the Association for Computational Linguistics, D. Jurafsky, J. Chai, N. Schluter, and J. Tetreault (Eds.), Online, pp. 3575–3585. External Links: Link, Document Cited by: §2.
  • P. Koehn, H. Hoang, A. Birch, C. Callison-Burch, M. Federico, N. Bertoldi, B. Cowan, W. Shen, C. Moran, R. Zens, C. Dyer, O. Bojar, A. Constantin, and E. Herbst (2007) Moses: open source toolkit for statistical machine translation. In Proceedings of the 45th Annual Meeting of the Association for Computational Linguistics Companion Volume Proceedings of the Demo and Poster Sessions, S. Ananiadou (Ed.), Prague, Czech Republic, pp. 177–180. External Links: Link Cited by: §3.1.
  • G. Kuwanto, C. Agarwal, G. I. Winata, and D. T. Wijaya (2024) Linguistics theory meets LLM: code-switched text generation via equivalence constrained large language models. External Links: 2410.22660, Link Cited by: §2, §6.
  • M. Li, W. Shi, A. Pagnoni, P. West, and A. Holtzman (2024) Predicting vs. acting: a trade-off between world modeling & agent modeling. ArXiv abs/2407.02446. External Links: Link Cited by: §5.1.
  • X. V. Lin, T. Mihaylov, M. Artetxe, T. Wang, S. Chen, D. Simig, M. Ott, N. Goyal, S. Bhosale, J. Du, R. Pasunuru, S. Shleifer, P. S. Koura, V. Chaudhary, B. O’Horo, J. Wang, L. Zettlemoyer, Z. Kozareva, M. Diab, V. Stoyanov, and X. Li (2022) Few-shot learning with multilingual generative language models. In Proceedings of the 2022 Conference on Empirical Methods in Natural Language Processing, Y. Goldberg, Z. Kozareva, and Y. Zhang (Eds.), Abu Dhabi, United Arab Emirates, pp. 9019–9052. External Links: Link, Document Cited by: §1.
  • G. Molina, F. AlGhamdi, M. Ghoneim, A. Hawwari, N. Rey-Villamizar, M. Diab, and T. Solorio (2016) Overview for the second shared task on language identification in code-switched data. In Proceedings of the Second Workshop on Computational Approaches to Code Switching, M. Diab, P. Fung, M. Ghoneim, J. Hirschberg, and T. Solorio (Eds.), Austin, Texas, pp. 40–49. External Links: Link, Document Cited by: §2.
  • NLLB Team, M. R. Costa-jussà, J. Cross, O. Çelebi, M. Elbayad, K. Heafield, K. Heffernan, E. Kalbassi, J. Lam, D. Licht, J. Maillard, A. Sun, S. Wang, G. Wenzek, A. Youngblood, B. Akula, L. Barrault, G. M. Gonzalez, P. Hansanti, J. Hoffman, S. Jarrett, K. R. Sadagopan, D. Rowe, S. L. Spruit, C. Tran, P. Y. Andrews, N. F. Ayan, S. Bhosale, S. Edunov, A. Fan, C. Gao, V. Goswami, F. Guzmán, P. Koehn, A. Mourachko, C. Ropers, S. Saleem, H. Schwenk, and J. Wang (2022) No language left behind: scaling human-centered machine translation. ArXiv abs/2207.04672. External Links: Link Cited by: §4.
  • K. Papineni, S. Roukos, T. Ward, and W. Zhu (2002) Bleu: a method for automatic evaluation of machine translation. In Proceedings of the 40th Annual Meeting of the Association for Computational Linguistics, P. Isabelle, E. Charniak, and D. Lin (Eds.), Philadelphia, Pennsylvania, USA, pp. 311–318. External Links: Link, Document Cited by: §6.1.
  • P. Patwa, G. Aguilar, S. Kar, S. Pandey, S. PYKL, B. Gambäck, T. Chakraborty, T. Solorio, and A. Das (2020) SemEval-2020 task 9: overview of sentiment analysis of code-mixed tweets. In Proceedings of the Fourteenth Workshop on Semantic Evaluation, A. Herbelot, X. Zhu, A. Palmer, N. Schneider, J. May, and E. Shutova (Eds.), Barcelona (online), pp. 774–790. External Links: Link, Document Cited by: §2.
  • S. Poplack (1980) Sometimes I’ll start a sentence in Spanish y termino en español: toward a typology of code-switching 1. Linguistics 18, pp. 581–618. External Links: Document Cited by: §1, §2.
  • M. Popović (2015) ChrF: character n-gram F-score for automatic MT evaluation. In Proceedings of the Tenth Workshop on Statistical Machine Translation, O. Bojar, R. Chatterjee, C. Federmann, B. Haddow, C. Hokamp, M. Huck, V. Logacheva, and P. Pecina (Eds.), Lisbon, Portugal, pp. 392–395. External Links: Link, Document Cited by: §6.1.
  • T. Potter and Z. Yuan (2024) LLM-based code-switched text generation for grammatical error correction. In Proceedings of the 2024 Conference on Empirical Methods in Natural Language Processing, Y. Al-Onaizan, M. Bansal, and Y. Chen (Eds.), Miami, Florida, USA, pp. 16957–16965. External Links: Link, Document Cited by: §2.
  • A. Pratapa, G. Bhat, M. Choudhury, S. Sitaram, S. Dandapat, and K. Bali (2018) Language modeling for code-mixing: the role of linguistic theory based synthetic data. In Proceedings of the 56th Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), I. Gurevych and Y. Miyao (Eds.), Melbourne, Australia, pp. 1543–1553. External Links: Link, Document Cited by: §2.
  • A. B. Sai, A. K. Mohankumar, and M. M. Khapra (2022) A survey of evaluation metrics used for nlg systems. ACM Comput. Surv. 55 (2). External Links: ISSN 0360-0300, Link, Document Cited by: §6.3.
  • E. Sarkisov (2021) Interlingual interference as a linguistic and cultural characteristic of the current online communication. Russian Journal of Bilingualism Studies 3, pp. 16–21. Cited by: §1.
  • T. Solorio, E. Blair, S. Maharjan, S. Bethard, M. Diab, M. Ghoneim, A. Hawwari, F. AlGhamdi, J. Hirschberg, A. Chang, and P. Fung (2014) Overview for the first shared task on language identification in code-switched data. In Proceedings of the First Workshop on Computational Approaches to Code Switching, M. Diab, J. Hirschberg, P. Fung, and T. Solorio (Eds.), Doha, Qatar, pp. 62–72. External Links: Link, Document Cited by: §2.
  • V. Srivastava and M. Singh (2021) Challenges and limitations with the metrics measuring the complexity of code-mixed text. In Proceedings of the Fifth Workshop on Computational Approaches to Linguistic Code-Switching, T. Solorio, S. Chen, A. W. Black, M. Diab, S. Sitaram, V. Soto, E. Yilmaz, and A. Srinivasan (Eds.), Online, pp. 6–14. External Links: Link, Document Cited by: §1, §6.
  • I. Tarunesh, S. Kumar, and P. Jyothi (2021) From machine translation to code-switching: generating high-quality code-switched text. In Proceedings of the 59th Annual Meeting of the Association for Computational Linguistics and the 11th International Joint Conference on Natural Language Processing (Volume 1: Long Papers), C. Zong, F. Xia, W. Li, and R. Navigli (Eds.), Online, pp. 3154–3169. External Links: Link, Document Cited by: §2.
  • M. Terblanche, K. Olaleye, and V. Marivate (2024) Prompting towards alleviating code-switched data scarcity in under-resourced languages with GPT as a pivot. In Proceedings of the 3rd Annual Meeting of the Special Interest Group on Under-resourced Languages @ LREC-COLING 2024, M. Melero, S. Sakti, and C. Soria (Eds.), Torino, Italia, pp. 272–282. External Links: Link Cited by: §2.
  • G. R. Tucker (2001) A global perspective on bilingualism and bilingual education. Georgetown University Round table on Languages and Linguistics 1999, pp. . Cited by: §1.
  • J. Wang, Y. Liang, F. Meng, Z. Sun, H. Shi, Z. Li, J. Xu, J. Qu, and J. Zhou (2023) Is ChatGPT a good NLG evaluator? a preliminary study. In Proceedings of the 4th New Frontiers in Summarization Workshop, Y. Dong, W. Xiao, L. Wang, F. Liu, and G. Carenini (Eds.), Singapore, pp. 1–11. External Links: Link, Document Cited by: §6.3, §6.
  • P. West and C. Potts (2025) Base models beat aligned models at randomness and creativity. In Second Conference on Language Modeling, External Links: Link Cited by: §5.1.
  • G. Winata, A. F. Aji, Z. X. Yong, and T. Solorio (2023) The decades progress on code-switching research in NLP: a systematic survey on trends and challenges. In Findings of the Association for Computational Linguistics: ACL 2023, A. Rogers, J. Boyd-Graber, and N. Okazaki (Eds.), Toronto, Canada, pp. 2936–2978. External Links: Link, Document Cited by: §2.
  • G. I. Winata, S. Cahyawijaya, Z. Liu, Z. Lin, A. Madotto, and P. Fung (2021) Are multilingual models effective in code-switching?. In Proceedings of the Fifth Workshop on Computational Approaches to Linguistic Code-Switching, T. Solorio, S. Chen, A. W. Black, M. Diab, S. Sitaram, V. Soto, E. Yilmaz, and A. Srinivasan (Eds.), Online, pp. 142–153. External Links: Link, Document Cited by: §1.
  • T. Wolf, L. Debut, V. Sanh, J. Chaumond, C. Delangue, A. Moi, P. Cistac, T. Rault, R. Louf, M. Funtowicz, J. Davison, S. Shleifer, P. von Platen, C. Ma, Y. Jernite, J. Plu, C. Xu, T. L. Scao, S. Gugger, M. Drame, Q. Lhoest, and A. M. Rush (2020) Transformers: state-of-the-art natural language processing. In Proceedings of the 2020 Conference on Empirical Methods in Natural Language Processing: System Demonstrations, Online, pp. 38–45. External Links: Link Cited by: §4.
  • S. Xu and M. F. Lorber (2014) Interrater agreement statistics with skewed data: evaluation of alternatives to cohen’s kappa. Journal of consulting and clinical psychology 82 (6), pp. 1219–1227. External Links: Link, Document Cited by: §3.2.1.
  • Z. X. Yong, R. Zhang, J. Forde, S. Wang, A. Subramonian, H. Lovenia, S. Cahyawijaya, G. Winata, L. Sutawika, J. C. B. Cruz, Y. L. Tan, L. Phan, L. Phan, R. Garcia, T. Solorio, and A. F. Aji (2023) Prompting multilingual large language models to generate code-mixed texts: the case of south East Asian languages. In Proceedings of the 6th Workshop on Computational Approaches to Linguistic Code-Switching, G. Winata, S. Kar, M. Zhukova, T. Solorio, M. Diab, S. Sitaram, M. Choudhury, and K. Bali (Eds.), Singapore, pp. 43–63. External Links: Link Cited by: §2.
  • T. Zhang, V. Kishore, F. Wu, K. Q. Weinberger, and Y. Artzi (2020) BERTScore: evaluating text generation with BERT. ArXiv abs/1904.09675. External Links: 1904.09675, Link Cited by: §6.1.
  • W. Zhu, H. Liu, Q. Dong, J. Xu, S. Huang, L. Kong, J. Chen, and L. Li (2024) Multilingual machine translation with large language models: empirical results and analysis. In Findings of the Association for Computational Linguistics: NAACL 2024, K. Duh, H. Gomez, and S. Bethard (Eds.), Mexico City, Mexico, pp. 2765–2781. External Links: Link, Document Cited by: §4.